OpenStudy (anonymous):

find the real solution x^-2

OpenStudy (anonymous):

\[x \neq 0\]

OpenStudy (anonymous):

\[x ^{-2}+9x ^{-1}+18=0\]

OpenStudy (anonymous):

that is the formula

OpenStudy (anonymous):

set x^-1 as u which makes x^-2 to u^2 and solve as a 2n degree equation

OpenStudy (anonymous):

Just swing the negative exponents in the denominator.

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

?

OpenStudy (anonymous):

u^2+9u+18=0 And then the answer to x should be 1/x1 and 1/x2

OpenStudy (anonymous):

i mean u1 and u2, to get the values of x1 and x2

OpenStudy (anonymous):

since x^-1 = 1/x and x^-2 = 1/x^2

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{2} }+\frac{ 9 }{ x }+18=0\]

OpenStudy (anonymous):

okay so what would be the answer?